OpenStudy (anonymous):

Which of the following best defines 4 to the power of 2 over 3? Square root of 16 Cube root of 4 Square root of 4 Cube root of 16

OpenStudy (anonymous):

\[4^{\frac{2}{3}}=\sqrt[3]{4^2}\]

OpenStudy (anonymous):

since \(4^2=16\) it is also \(\sqrt[3]{16}\)

OpenStudy (anonymous):

\[4^{2/3}=(4^{1/3})^{2}=(\sqrt{2})^{2}=2\]

OpenStudy (anonymous):

Therefore, 2 is the square root of 4.

OpenStudy (anonymous):

nooooo because \[\sqrt[3]{4}^2 \] so it would be 3sqr16

OpenStudy (wwwhhhaaattt?):

@satellite73 Which answer is correct?

OpenStudy (domebotnos):

yeah what is the answer choice???

OpenStudy (anonymous):

B